Use the following information to answer the next three questions.

QUESTION 5

As of today, the spot exchange rate is £1.25/$. The U.S. interest rate is 7% and the interest rate in the euro zone is 10%. What is the one-year forward rate (in terms of a direct quote from the US view) that should prevail according to IRP? Round intermediate steps and your final answer to four decimals. Assume the US is the domestic country. Do not use currency symbols or words when entering your response.

QUESTION 6

Suppose that the one year forward rate is $.75/£ . Find the profit (in terms of percentage returns) you could earn via covered interest arbitrage. Assume the US is the home country. Round intermediate steps to four decimals.

.1033

.0387

.0669

.0433

QUESTION 7

Which of the following could occur to eliminate the arbitrage opportunity?

An increase in the spot rate for pounds in terms of dollars.

A decrease in the forward rate for pounds in terms of dollars.

A decrease in the spot rate for pounds in terms of dollars.

None of the above
Answer #1

5.As per Interest Rate parity, forward rate = Spot Rate (1+Interest Rate Pound)/(1+Interest Rate Dollar)

= 1.25(1+10%)/(1+7%)

= Pound 1.2850/Dollar

One year forward rate = $0.75/Pound or 1.333 Pound/Dollar

Borrow 1.25 pound

Convert into USD and invest for one year and get 1(1.07) = $1.07

Convert back into pound = 1.07*1.3333 = Pound 1.4266

Cost of pound after interest = 1.25(1.1) = Pound 1.375

% return = (1.4266-1.375)/1.375 = 0.0387 (approx.)

A decrease in the forward rate for pounds in terms of dollars.

Since pound is overvalues in forward market